An urn contains 9 pink and 7 black balls. Four balls are randomly drawn from the urn in succession, with replacement. That is, after each draw, the selected ball is returned to the urn. What is the probability that all balls drawn from the urn are pink? Round your answer to three decimal places.
Answer by Edwin McCravy(20064)

An urn contains 9 pink and 7 black balls. Four balls are randomly drawn from the
urn in succession, with replacement. That is, after each draw, the selected ball
is returned to the urn. What is the probability that all balls drawn from the
urn are pink? Round your answer to three decimal places.
There are 9+7=16 balls in the urn each time

Therefore the probability of drawing a pink ball each time is 9/16

P(pink 1st AND pink 2nd AND pink 3rd AND pink 4th) =

P(pink 1st)*P(pink 2nd)*P(pink 3rd)*P(pink 4th) =

(9/16)*(9/16)*(9/16)*(9/16) =

(9/16)^4

9^4/16^4

6561/65536

About 0.100112915 or about 0.1

Edwin